OpenStudy (anonymous):

K is the midpoint of line AB. If line AK=3x+1 and line KB=1/3x+9, find the length of AK.

OpenStudy (anonymous):

ok A.......K.......B well AK and KB are equal so this has to be true: 3x+1=1/3x+9 now we have to get x by itself.....subtract 1 from both sides 3x=1/3x+8 now subtract 1/3x from both sides 3x-1/3x=8 you are left with...2.667x=8 divide both sides by 2.667 and you get x=2.999 Rounded up is 3=x now plug x(3) into either one of the segments and multiply your result by 2 so....3(3) +1=10*2=20! 20 units is the answer hope this helps =]

OpenStudy (anonymous):

Thank you :3